OpenStudy (anonymous):

fr33 or kwe help me on that integral questions : ill post it to down

OpenStudy (anonymous):

\[\int\limits_{}\sqrt{tanx}*dx\]

OpenStudy (anonymous):

well I don't know this, I suck at calculus :(

OpenStudy (anonymous):

lol...

OpenStudy (anonymous):

\[u=sinx/cosx\] then \[du=\sin2-\cos2x*dr\]

OpenStudy (anonymous):

then it will be that \[\int\limits_{}\sqrt{sinx/cosx}*\sin2x-\cos2x*dx\]

OpenStudy (anonymous):

where u got the 2? from the index of the radical?

OpenStudy (anonymous):

its 2x not 2 sorr

OpenStudy (anonymous):

du=sin2x−cos2x∗dr

OpenStudy (anonymous):

yup I am lost :p

OpenStudy (anonymous):

:P